# Attention-Seeker: Dynamic Self-Attention Scoring for Unsupervised Keyphrase Extraction

### v1

@@ 1 Introduction
Unsupervised keyphrase extraction methods can be divided into five categories based on the type of information they extract. The first two categories include traditional approaches: statistical methods (Spärck Jones, 2004; Campos et al., 2020), which use in-corpus statistical information, and graph-based methods (Mihalcea and Tarau, 2004; Wan and Xiao, 2008; Bougouin et al., 2013; Florescu and Caragea, 2017), which leverage co-occurrence patterns. The remaining three categories consist of more recent works that use Pretrained Language Models (PLMs) to extract semantic information. Embedding-based methods Bennani-Smires et al. (2018); Sun et al. (2020); Ding and Luo (2021); Zhang et al. (2022); Liang et al. (2021) analyze similarities between documents and phrases within the PLM’s embedding space. Prompt-based methods (Kong et al., 2023) use the PLM’s decoder logits to estimate the probability of generating a candidate phrase when a prompt is included in the document (e.g., this book talks about [candidate]). Finally, self-attention-based methods (Kang and Shin, 2023) examine the PLMs’ Self-Attention Map (SAM) to identify which candidate phrases receive the most attention across all document tokens.
@@
In this context, we propose a method that automatically adapts to the characteristics of the input text, thereby eliminating the need for manual parameter tuning. Our approach extends SAMRank (Kang and Shin, 2023), introducing a module that selects the most relevant SAMs to score candidate phrases effectively. Since a direct evaluation of SAMs using labeled keyphrases is impractical, we estimate SAMs’ relevance by defining the characteristics of an optimal SAM. Specifically, Attention-Seeker defines a binary vector that assigns ones to candidate phrases and zeros to all other tokens. It then computes the similarity between this vector and SAMs’ attention vectors (the rows of the self-attention maps), and use the average similarity score to determine the relevance of each SAM. For short documents, Attention-Seeker further evaluates the relevance of individual attention vectors, while for long documents, it considers the relevance of the document segments, refining the attention score estimate for the candidate phrases.
To the best of our knowledge, Attention-Seeker is the first method that automatically adapts its keyphrase extraction process to the specific characteristics of the input text. This novel approach allows Attention-Seeker to achieve SOTA-level performance on four benchmark datasets without requiring any parameter tuning. Our ablation study shows that each of the estimated relevance scores in our method contributes positively to the improvement of a non-parametric SAMRank. Although our proposed relevance scoring mechanism is simple, its effectiveness suggests that further exploration in this direction would be beneficial.
@@ 2.2 Self-Attention Map
In the transformer model proposed by Vaswani et al. (2017), and in subsequent PLMs, each layer contains multiple heads. As a result, each SAM can focus on different aspects of text processing (multiple heads) at different syntactic levels (multiple layers). Clark et al. (2019) examined the characteristics of different SAMs in the BERT model (Devlin et al., 2019) and identified several types: SAMs that specialize in attending to previous or next tokens, SAMs that focus on separator tokens (e.g., special tokens, periods, and commas), and SAMs that attend to a large number of words (uniform attention). In addition, they identified specific SAMs that preserve syntactic information, suggesting that certain SAMs may also specialize in attending to key topics within the text (Kang and Shin, 2023).
@@ 3 Methodology
In this section, we introduce our proposed method, Attention-Seeker. It consists of four main steps: (1) Generation of candidate phrases using Part-of-Speech (POS) sequences. (2) Extraction of Self-Attention Maps (SAMs). (3) Estimation of a vector of attention scores from the most relevant SAMs. (4) Scoring the importance of each candidate phrase based on these attention scores. Our method extends SAMRank (Kang and Shin, 2023) by proposing a new approach for estimating the vector of attention scores from all the SAMs (step 3), rather than relying on the manual selection of a single SAM (parameter tuning).
3.1 Candidate Generation
We follow a well-established approach for generating candidate phrases (Kang and Shin, 2023; Ding and Luo, 2021; Bennani-Smires et al., 2018). First, we tokenize and POS tag all words in the document by using the Standford CoreNLP tool [^1] https://stanfordnlp.github.io/CoreNLP/. Next, we extract noun phrases (tagged as “NN”, “NNS”, “NNP”, “NNPS”, and “JJ”) using the NLTK’s RegexpParser [^2] https://github.com/nltk. These phrases are then defined as the document’s keyphrase candidates.
3.2 Extraction of the Self-Attention Maps
We extract the SAMs from each layer and head of the Large Language Model LLAMA 3-8B (AI@Meta, 2024) using the Huggingface library [^3] https://github.com/huggingface/transformers. Similar to SAMRank (Kang and Shin, 2023), we handle the extraction of SAMs differently depending on the document length. For short documents, we input the entire text directly into the LLM. For long documents, we first input the document’s abstract into the LLM, then divide the remaining text into segments with a consistent number of tokens, and process each segment separately.
@@ 3.3 Attention Scores Estimation: Short Documents
As shown in Figure 1, the "Attention Score Estimation" module first estimates the relevance of SAMs and their internal attention vectors. These relevance values are then used to compute a weighted average of the attention vectors across all SAMs. Accordingly, the final vector of attention scores $B$ is determined from the most relevant parts of the most relevant SAMs.
To estimate the relevance of each attention vector, we define a hypothesized vector $H$ that characterizes the desired features of a relevant SAM vector. Our initial hypothesis is that SAMs specialized in capturing keyphrases pay attention to tokens belonging to candidate phrases (Equation  2). For instance, SAMs that specialize in attending to token positions or separator tokens are not relevant to our task (Clark et al., 2019).
@@
Given a SAM $A^{lh}$, where $l$ is the current layer and $h$ is the current head, the relevance $S^{lh}_{i}$ of each attention vector $A^{lh}_{i}$ is calculated as the dot product between $H$ and $A^{lh}_{i}$. This calculation is formally expressed as the matrix multiplication shown in Equation  3, where $S^{lh}$ is the vector of $S^{lh}_{i}$ values.
@@
We further set $S^{lh}_{i}$ to 00 for tokens $i$ that do not belong to candidate phrases. We consider their corresponding attention vectors may introduce noise, as they are mapped to the contextualized embeddings of non-candidate tokens in subsequent layers. For example, attention vectors for periods tend to be uniform, helping to create a rich contextual embedding in preparation for a new context.
To determine the relevance $R^{lh}$ of each SAM, we calculate the average relevance score of its individual attention vectors (Equation  4). The final attention vector $B$ is then computed as a weighted average of all attention vectors across all SAMs, with the weights determined by their respective relevance scores (Equation  5).
@@
As shown in Figure  1, the obtained vector $B$ is converted into a new hypothesized vector $H$ by setting to 00 the values $B_{i}$ corresponding to non-candidate tokens. This results in a refined vector of attention scores $B$, now based on the hypothesis that SAMs should focus on tokens belonging to the most important candidate phrases.
@@ 3.4 Attention Scores Estimation: Long Documents
As described earlier, long documents are segmented into abstract and equally sized segments. Following this segmentation, we estimate attention scores for each segment using the same approach as for short documents (Figure  2). The main difference from our previous method is that relevance scores for attention vectors are not included in the final calculation. This adjustment helps avoid the complexities of applying context-specific relevance scores across multiple, heterogeneous contexts, which could reduce effectiveness without a robust normalization strategy.
Thus, for the segment $s$, the relevance $R^{lh}$ of each of its SAMs is defined as the average of the relevance scores of its corresponding attention vectors (Equation  6). We then apply $l_{1}$-normalization to this vector (Equation  7) to compute a weighted average $SAM^{s}$ of all SAMs from the segment $s$ (Equation  8). This normalization ensures that $SAM^{s}$ preserves the standard properties of a SAM (e.g., the attention values in each row sum to 1). Finally, the attention score vector $B^{s}$ for the segment $s$ is obtained by summing all rows in the average SAM (Equation  9).
@@
In addition, we define the relevance $T^{s}$ of each segment. To this end, we define a new hypothesized vector $H^{s}$, whose values correspond to the attention scores obtained in the abstract ($B^{0}$) for tokens that belong to the candidate phrases of the abstract, and 0 for all other tokens (Equation  10). Thus, the relevance score $T^{s}$ of the segment $s$ is calculated as the dot product of its attention score vector $B^{s}$ and this newly hypothesized vector (Equation  11).
@@ 4.2 Baselines
Our baselines include statistics-based methods: TF-IDF (Spärck Jones, 2004), YAKE (Campos et al., 2020); graph-based methods: TextRank (Mihalcea and Tarau, 2004), SingleRank (Wan and Xiao, 2008), TopicRank (Bougouin et al., 2013), PositionRank (Florescu and Caragea, 2017); embedding-based methods: EmbedRank (Bennani-Smires et al., 2018), SIFRank (Sun et al., 2020), AttentionRank (Ding and Luo, 2021), MDERank (Zhang et al., 2022), JointGL (Liang et al., 2021); prompt-based methods: PromptRank (Kong et al., 2023), and self-attention-based models: SAMRank (Kang and Shin, 2023). To ensure a fair comparison, we implemented SAMRank using the average of all SAMs (a non-parametric method) extracted with the LLAMA 3-8B model. The optimized results of SAMRank are detailed in Appendix  A. Additionally, we omitted the proportional score, as it led to lower performance in this setting (Appendix  B). Consequently, the reported results of SAMRank represent the most important baseline for our model, as they correspond to the same method when all relevance scores are assigned equal values.
@@ 5.1 Short documents
Attention-Seeker integrates several modules designed to improve its overall performance. Specifically, it employs a relevance score for each SAM’s attention vector ($S^{lh}$), a relevance score for each SAM ($R^{lh}$), and a filtering step $f$ applied to non-candidate tokens of $S^{lh}$ (Section  3.3). To evaluate their contributions, we applied these modules separately. We also tested configurations using both scores $S^{lh}$ and $R^{lh}$, with and without the filtering step. Table  3 presents the performance of Attention-Seeker in all these cases. In the "Base" configuration, all attention vectors and SAMs are assigned equal relevance, and no filtering is applied. The following configurations progressively incorporate the relevance scores ($S^{lh}$ and $R^{lh}$) and the filtering step ($f$) into the base method ($B$). The final version of Attention-Seeker corresponds to "$B+S^{lh}+R^{lh}+f$" applied twice (Figure  1). Performance on the SemEval2010 and Krapivin datasets is evaluated using only the abstracts of the papers (short documents).
@@ 5.2 Long documents
Table  4 shows that our two proposed relevance scores contribute positively to the performance of Attention-Seeker, with the relevance of segments providing the most significant improvements. In addition, refining the hypothesized vector $H^{s}$ further improves the performance of the segment scores ($T^{s}>T^{s}_{b}$). Despite these advances, our segmentation approach introduces some performance degradation compared to the base model, likely due to the difference in length between the abstract and other segments. While the overall improvement achieved by our method outweighs the initial degradation, further research may consider exploring new methods for estimating segment relevance without compromising the performance of the base model.
@@ 5.3 Relevance of SAMs
Attention-Seeker’s score $R^{lh}$ identifies the SAMs most relevant for keyphrase extraction, eliminating the need for labeled data to select an optimal SAM. Our analysis of the $R^{lh}$ scores across samples from four datasets (Inspec, SemEval2017, SemEval2010, and Krapivin) reveals that SAMs from intermediate layers (9 to 15 out of 32), along with a few SAMs from the first and last layer contribute significantly to the keyphrase extraction task (see Appendix  C).

### v2

@@ 1 Introduction
Unsupervised keyphrase extraction methods can be divided into five categories based on the type of information they extract. The first two categories include traditional approaches: statistical methods (Spärck Jones, 2004; Campos et al., 2020), which use in-corpus statistical information, and graph-based methods (Mihalcea and Tarau, 2004; Wan and Xiao, 2008; Bougouin et al., 2013; Florescu and Caragea, 2017), which leverage words co-occurrence patterns. The remaining three categories consist of more recent works that use Pretrained Language Models (PLMs) to extract semantic information. Embedding-based methods Bennani-Smires et al. (2018); Sun et al. (2020); Ding and Luo (2021); Zhang et al. (2022); Liang et al. (2021) analyze similarities between documents and phrases within the PLM’s embedding space. Prompt-based methods (Kong et al., 2023) use the PLM’s decoder logits to estimate the probability of generating a candidate phrase when a prompt is included in the document (e.g., this book talks about [candidate]). Finally, self-attention-based methods (Kang and Shin, 2023) examine the PLMs’ Self-Attention Map (SAM) to identify which candidate phrases receive the most attention across all document tokens.
@@
In this context, we propose a method that automatically adapts to the characteristics of the input text, eliminating the need for manual parameter tuning. Our approach extends SAMRank (Kang and Shin, 2023) by introducing a module that selects the most relevant SAMs to effectively score candidate phrases. The relevance of SAMs is estimated by defining the characteristics of an optimal SAM. Specifically, Attention-Seeker employs a binary vector that assigns ones to candidate phrases and zeros to all other tokens. It then computes the similarity between this vector and the attention vectors (rows) of the SAMs, using the average similarity score to determine each SAM’s relevance. Furthermore, Attention-Seeker evaluates the relevance of different parts of the text: for short documents, it scores individual attention vectors, whereas for long documents, it scores document segments.
To the best of our knowledge, Attention-Seeker is the first method to automatically adapt its keyphrase extraction process to the specific characteristics of the input text. This novel approach allows Attention-Seeker to achieve SOTA-level performance on four benchmark datasets without requiring parameter tuning. Our ablation study shows that each of the Attention-Seeker’s scoring modules positively contributes to the final performance. While our proposed relevance scoring mechanism is simple, its effectiveness highlights the potential for further exploration in this direction.
@@ 2.2 Self-Attention Map
In the transformer model proposed by Vaswani et al. (2017), and in subsequent PLMs, each layer contains multiple heads. As a result, each SAM can focus on different aspects of text processing (multiple heads) at different syntactic levels (multiple layers). Clark et al. (2019) examined the characteristics of different SAMs in the BERT model (Devlin et al., 2019) and identified several types: SAMs that specialize in attending to previous or next tokens, SAMs that focus on separator tokens (e.g., special tokens, periods, and commas), and SAMs that attend to a large number of words (uniform attention). Furthermore, they identified specific SAMs that preserve syntactic information, suggesting that some SAMs might specialize in attending to keyphrases within the text. Supporting this hypothesis, Kang and Shin (2023) demonstrated that attention scores from a manually selected SAM can be leveraged to achieve SOTA performance in keyphrase extraction tasks. However, identifying and selecting SAMs specialized in attending to keyphrases remains an open challenge in this approach.
@@ 3 Methodology
In this section, we introduce our proposed method, Attention-Seeker. It consists of four main steps: (1) Generation of candidate phrases using Part-of-Speech (POS) sequences. (2) Extraction of Self-Attention Maps (SAMs). (3) Estimation of a vector of attention scores from the most relevant SAMs. (4) Scoring the importance of each candidate phrase based on these attention scores. The main contribution of our method lies in step 3, where we propose a novel approach to estimate the vector of attention scores by identifying and exploiting information from the most relevant SAMs.
3.1 Candidate Generation
We follow a well-established approach for generating candidate phrases (Kang and Shin, 2023; Ding and Luo, 2021; Bennani-Smires et al., 2018). First, we tokenize and POS tag all words in the document by using the Stanford CoreNLP tool [^1] https://stanfordnlp.github.io/CoreNLP/. Next, we extract noun phrases (tagged as “NN”, “NNS”, “NNP”, “NNPS”, and “JJ”) using the NLTK’s RegexpParser [^2] https://github.com/nltk. These phrases are then defined as the document’s keyphrase candidates.
Figure 1: The effect of different hypothesis vectors H in an LLM’s distribution of attention scores.
3.2 Extraction of the Self-Attention Maps
We extract the SAMs from each layer and head of the Large Language Model LLAMA 3-8B (AI@Meta, 2024) using the Huggingface library [^3] https://github.com/huggingface/transformers. Similar to SAMRank (Kang and Shin, 2023), we handle the extraction of SAMs differently depending on the document length. For short documents, we input the entire text directly into the LLM. For long documents, we first input the document’s abstract into the LLM, then divide the remaining text into equally sized segments (same number of tokens) and process each segment independently with the LLM.
3.3 SAMs’ Relevance Scoring: Hypothesis Engineering
All Self-Attention Maps (SAMs) in a Large Language Model (LLM) capture distinct attention patterns over tokens, and their distribution of attention scores serves as a unique fingerprint. We propose to use these fingerprints to measure the relevance of each SAM within the LLM. Given a SAM of size $n\times n$, our method requires the design of a hypothesis vector $H$ of length $n$ that encapsulates the desired properties of the attention distribution for a given task. Then, we calculate a similarity score between the SAM and the vector $H$ that quantifies the relevance of the SAM to the given task. We use this score as a weight to combine the attention distributions of all SAMs into a single output distribution for the LLM. This averaged distribution aggregates information across the different contexts captured by all layers and heads, prioritizing the contributions of the most relevant SAMs.
Figure 1 illustrates the effect of a hypothesis vector $H$ on the output distribution of attention scores for the same input text. The first case represents the default output distribution of the LLM, where all SAMs are considered equally relevant. In this scenario, $H$ is defined as a uniform vector with all elements set to $1$, resulting in equal similarity scores for all SAMs. In the second case, $H$ is defined as a binary vector where only the elements corresponding to “dog” are set to $1$. This configuration assigns greater relevance to SAMs that focus on syntactic relationships involving the noun “dog” and facilitates the retrieval of information about “dog”. A simple post-processing step, which masks the attention scores of “dog” from the final distribution, allows the extraction of associated words such as “playful” and “chased”.
The third and fourth cases demonstrate the use of combined hypothesis vectors. Both cases include a hypothesis vector designed to prioritize SAMs that capture syntactic information related to verbs (assigning $1$ to tokens corresponding to verbs and 00 to others). In the third case, this verb-focused vector is combined with the hypothesis vector for “dog”, while in the fourth case, it is combined with the vector for “tail”. As shown in Figure 1, the resulting distributions differ significantly, highlighting the verb most related to “dog” in the third case and the verb most related to “tail” in the fourth case. These examples illustrate how task-specific hypothesis vectors can guide the extraction of syntactically relevant information.
Although the examples shown in Figure 1 use simple hypothesis vectors, future works could explore advanced methodologies to design optimal vectors tailored to specific tasks. In this study, we propose two simple designs for keyphrase extraction: one for short and another for long documents. Both designs start from a vector $H$ that focuses on the keyphrase candidates. This approach prioritizes SAMs that model the inter-relationships of these candidates. Among the relevant SAMs, we can include those capturing long-term dependencies, object-direct relationships, coreferences, and keyphrases.
@@ 3.4 Attention Scores Estimation: Short Documents
Our proposed method for short documents is summarized in Figure 2. Here, the "Attention Score Estimation" module first estimates the relevance of SAMs ($R^{lh}$) and their internal attention vectors ($S^{lh}_{i}$). These relevance values are then used to compute a weighted average of the attention vectors across all SAMs. Accordingly, the final vector of attention scores $B$ is determined from the most relevant parts of the most relevant SAMs.
We define the hypothesis vector $H$ as in Equation 2. This vector prioritizes SAMs that pay attention to potential keyphrases (candidates), while de-emphasizing those that primarily attend to token positions or separator tokens (Clark et al., 2019).
@@
Given a SAM $A^{lh}$, where $l$ is the current layer and $h$ is the current head, the relevance of each attention vector $S^{lh}_{i}$ is calculated as the dot product between $H$ and $A^{lh}_{i}$. This calculation is formally expressed as the matrix multiplication shown in Equation 3, where $S^{lh}$ is the vector of $S^{lh}_{i}$ values.
@@
We further set $S^{lh}_{i}$ to 00 for tokens $i$ that do not belong to candidate phrases (post-processing masking). This is an empirical decision, and can be justified by the existence of uniformly distributed attention vectors (e.g., periods preserving contextual information for subsequent sentences) or randomly distributed attention vectors for "non-op" contexts (Clark et al., 2019).
As shown by Equation 4, the relevance of each SAM $R^{lh}$ is calculated as the average relevance score of its individual attention vectors.
@@
The final attention vector $B$ is computed as a weighted average of all attention vectors across all SAMs, with the weights determined by their respective relevance scores, as shown in Equation 5.
@@
As illustrated in Figure 2, the resulting vector $B$ is converted into a new hypothesis vector $H$ through post-processing masking, where $B_{i}$ values corresponding to non-candidate tokens are set to 0. Similar to the initial binary vector $H$, this updated vector prioritizes SAMs that focus on possible keyphrases. However, it further emphasizes SAMs that allocate more attention to the most important candidates, thereby favoring SAMs likely to specialize in attending to keyphrases.
@@ 3.5 Attention Scores Estimation: Long Documents
Long documents are segmented into an abstract and equally sized segments, as described in Section 3.2. These segments are processed independently as short documents (Section 3.4) excluding the relevance scores for attention vectors ($S^{lh}_{i}$) from the final calculation (Figure 3). This simplification avoids the complexities of applying context-dependent relevance scores across diverse segments, which could reduce effectiveness without a robust normalization strategy.
For each segment $s$, the relevance of each SAM ($R^{lh}$) is defined as the average of the relevance scores of its corresponding attention vectors (Equation  6).
@@
We then apply $l_{1}$-normalization to this vector (Equation 7) to compute a weighted average $SAM^{s}$ of all SAMs from the segment $s$ (Equation 8). This normalization ensures that $SAM^{s}$ preserves the standard properties of a SAM (e.g., the attention values in each row sum to 1).
@@
The attention score vector $B^{s}$ for the segment $s$ is obtained by summing all rows in the average SAM (Equation 9).
@@
To compute the relevance of each segment ($T^{s}$), we define a new hypothesis vector $H^{s}$ for each segment. This vector assigns the attention scores from the abstract ($B^{0}$) to tokens in the segment that are included in the candidate phrases of the abstract, while assigning 0 to all other tokens (Equation  10). This hypothesis vector prioritizes segments that focus more on phrases considered important by the abstract.
@@
The relevance score $T^{s}$ for each segment $s$ is finally calculated as the dot product between its attention score vector $B^{s}$ and the hypothesis vector $H^{s}$ (Equation  11).
@@ 4.2 Baselines
Our baselines include statistics-based methods: TF-IDF (Spärck Jones, 2004), YAKE (Campos et al., 2020); graph-based methods: TextRank (Mihalcea and Tarau, 2004), SingleRank (Wan and Xiao, 2008), TopicRank (Bougouin et al., 2013), PositionRank (Florescu and Caragea, 2017); embedding-based methods: EmbedRank (Bennani-Smires et al., 2018), SIFRank (Sun et al., 2020), AttentionRank (Ding and Luo, 2021), MDERank (Zhang et al., 2022), JointGL (Liang et al., 2021); prompt-based methods: PromptRank (Kong et al., 2023), and self-attention-based models: SAMRank (Kang and Shin, 2023). To ensure a fair comparison, we implemented SAMRank using the average of all SAMs (a non-parametric method) extracted with the LLAMA 3-8B model (BERT and GPT-2 results can be found on Appendix N). The optimized results of SAMRank are detailed in Appendix  A. Additionally, we omitted the proportional score, as it led to lower performance in this setting (Appendix  C). Consequently, the reported results of SAMRank represent the most important baseline for our model, as they correspond to the same method when all relevance scores are assigned equal values.
@@ 5.1 Short documents
Attention-Seeker integrates several modules designed to improve its overall performance. Specifically, it employs a relevance score for each SAM’s attention vector ($S^{lh}$), a relevance score for each SAM ($R^{lh}$), and a filtering step $f$ applied to non-candidate tokens of $S^{lh}$ (Section  3.4). To evaluate their contributions, we applied these modules separately. We also tested configurations using both scores $S^{lh}$ and $R^{lh}$, with and without the filtering step. Table  3 presents the performance of Attention-Seeker in all these cases. In the "Base" configuration, all attention vectors and SAMs are assigned equal relevance, and no filtering is applied. The following configurations progressively incorporate the relevance scores ($S^{lh}$ and $R^{lh}$) and the filtering step ($f$) into the base method ($B$). The final version of Attention-Seeker corresponds to "$B+S^{lh}+R^{lh}+f$" applied twice (Figure  2). Performance on the SemEval2010 and Krapivin datasets is evaluated using only the abstracts of the papers (short documents).
@@ 5.2 Long documents
Table  4 shows that our two proposed relevance scores contribute positively to the performance of Attention-Seeker, with the relevance of segments providing the most significant improvements. In addition, refining the hypothesis vector $H^{s}$ further improves the performance of the segment scores ($T^{s}>T^{s}_{b}$). Despite these advances, our segmentation approach introduces some performance degradation compared to the base model, likely due to the difference in length between the abstract and other segments. While the overall improvement achieved by our method outweighs the initial degradation, further research may consider exploring new methods for estimating segment relevance without compromising the performance of the base model.
@@ 5.3 Relevance of SAMs
Attention-Seeker’s score $R^{lh}$ identifies the SAMs most relevant for keyphrase extraction, eliminating the need for labeled data to select an optimal SAM. Our analysis of the $R^{lh}$ scores across samples from four datasets (Inspec, SemEval2017, SemEval2010, and Krapivin) reveals that SAMs from intermediate layers (9 to 15 out of 32), along with a few SAMs from the first and last layer contribute significantly to the keyphrase extraction task (see Appendix  D).